train a and train b are traveling in the same direction on parallel tracks. train a is traveling at 60 miles per hour and train b is traveling at 70 miles per hour. train a passes a station at 6:15 a.m. if train b passes the same station at 6:45 a.m, at what time will train b catch up to train a
Answer by mananth(16949)

Train A speed 60 mph
train B 70 mph.
Train A passes a station at 6:15 am.
If train B passes the same station at 6:45 am,
..
Speed train B - speed train A = catch up speed
70-60 = 10 mph.
..
time between trains passing the station = 1/2 hour
..
in 1/2 hour train A travels 60*1/2 = 30 miles.
...
catch up time = catch up distance / catch up speed
catch up time = 30/10= 3 hours
..
Train B will catch up with Train A at 6.45 + 3.00= 9:45 am
..
m.ananth@hotmail.ca